OpenStudy (anonymous):

Given the exponential equation 3^x=243, what is the logarithmic form of the equation in base 10?

OpenStudy (anonymous):

@phi

OpenStudy (phi):

take the log (base 10) of both sides

OpenStudy (anonymous):

How do I go about doing that?

OpenStudy (phi):

\[ \log(3^x)=\log(243) \]

OpenStudy (anonymous):

I'm sorry- I know that's pretty basic stuff, But its been over a month and I forgot most of it.

OpenStudy (anonymous):

Is adding the log before the parentheses all you have to do to take the log base 10 of both sides?

OpenStudy (phi):

"take the logarithm of x" means write log(x) in place of x. once we have done that, we might want to simplify. For example, we could use this rule: \[ \log(a^b) = b \log(a) \]

OpenStudy (anonymous):

ok- that would give us \[xlog(3)=\log(243)\]

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

@phi

OpenStudy (phi):

yes. and depending on the answer choices, you might want to write it as \[ x = \frac{\log(243)}{\log(3)} \]

OpenStudy (phi):

but they may want the form you wrote down.

OpenStudy (anonymous):

That's exactly how it's written, except they included the subscript of 10. Thank you for the help!